OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = 5 cos 5θ

OpenStudy (freckles):

keep in mind cos is an even function (that is cos(-t)=cos(t) ) you have r=5cos(5t) see if you have (r,-t) on the graph if you do then it is symmetric about the x-axis see if you have (-r,-t) on the graph if you do then it is symmetric about the y-axis see if you have (-r,t) on the graph if you do then it is symmetric about the origin

OpenStudy (anonymous):

is it symmetric on the y-axis? @freckles

OpenStudy (freckles):

how do you figure that?

OpenStudy (anonymous):

idk

OpenStudy (freckles):

here an example: let's look at r=sin(5t) sin is an odd function (that is sin(-t)=-sin(t)) so let's check x-axis symmetry? is (r,-t) on the curve? r=sin(5*-t) r=-sin(5t) as you see the equation r=-sin(5t) is not the same as r=sin(5t) so no we don't have x-axis symmetry so let's check y-axis symmetry? is (-r,-t) on the curve? -r=sin(5*-t) -r=-sin(5t) r=sin(5t) this is the same as what our equation r=sin(5t) so yes we have y-axis symmetry so let's check origin symmetry? is (-r,t) on the curve? -r=sin(5t) no r=-sin(5t) is not the same as r=sin(5t) so no we don't have origin symmetry all you have to do is use the guideline I gave you above

OpenStudy (freckles):

go through your check list the first thing on the check list is (r,-t) on the curve? plug in and see if

OpenStudy (freckles):

r = 5 cos 5t can you plug in (r,-t) into that equation?